(NOS. 9-10) PROBLEM 3. A silo has the form of a right circular cylinder 20m in diameter and 45m
high. The lateral surface of the silo is to be painted with two coats of elastomeric paint.

______________________ 9) What is the lateral surface area of the silo in m2?

______________________ 10) If the contract calls for two coats of paint and if it is estimated that
one gallon of paint can cover 30 square meters of surface in one coating, how many gallons of paint needed to cover the lateral surface area of the silo?
1
Expert's answer
2020-12-01T15:23:42-0500

9. Lateral surface area = 2πrh

Height, h = 45m

diameter, d = 20m

Radius, r = d/2 = 20/2 = 10m

Lateral surface area = 2*(22/7)*10*45

= 2828.57m²


10. 1 gallon = 30m²

x gallon = 2828.57m²

x = (2828.57*1)/30

x = 94.29 gallons

Therefore, 94.29 gallons = 2828.57m²

Since it requires two coats, the number of gallons needed will be 94.29 * 2 = 188.58 gallons.